show that cosx - cos^3 x over sinx is equal to sinxcosx


1
Expert's answer
2020-10-06T18:25:04-0400

"=cosx\/sinx - cos ^3x\/sinx" "=1\/tanx-cos^2 x\/tanx"

"=(1-cos^2x)\/tanx"

"=sin^2x\/tanx"

"=sin^2x\/(sinx\/cosx)"

"=sinxcosx"